OpenStudy (anonymous):

Can someone please help me with Solving Quadratic Equations with Complex Solutions????

OpenStudy (anonymous):

Post your equation

OpenStudy (anonymous):

Solve x2 + 2x + 9 = 0.

OpenStudy (anonymous):

@SithsAndGiggles

OpenStudy (anonymous):

You can use the quadratic formula directly: \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] with \(a=1,b=2,c=9\), then simplify the results. You can also complete the square: \[\begin{align*} x^2+2x+9&=x^2+2x+1+8\\ 0&=(x+1)^2+8 \end{align*}\] This makes it somewhat easier to isolate \(x\) if you're not comfortable with the previous formula (though you should definitely be comfortable with it, it's pretty important). A slightly harder way would be to factorize. This will give you the roots immediately. With a general equation, however, you might want to try a system of equations approach. We have \[ax^2+bx+c=x^2+2x+9\] and we want something of the form \(a(x-x_1)(x-x_2)\), where \(x_1\) and \(x_2\) are the roots/solutions. Expanding this, we get \[ax^2-ax_1x-ax_2x+ax_1x_2\] Matching up terms, we have \[\begin{align*}\color{red}ax^2+\color{blue}{(-ax_1-ax_2)}x+\color{green}{ax_1x_2}&=\color{red}{1}x^2+\color{blue}2x+\color{green}9\end{align*}\] so you have \[\begin{cases} a=1\\ -ax_1-ax_2=2\\ ax_1x_2=9 \end{cases}~~\iff~~ \begin{cases} -x_1-x_2=2\\ x_1x_2=9 \end{cases}\] and from here you solve for \(x_1\) and \(x_2\).